OpenStudy (anonymous):

when graphing, how do I know when to shade and when not to shade. my teacher told me to look at the signs but i still dont understand

OpenStudy (psymon):

Well, lets just pick something simple like x + y >= 2 You had when you can pick an x and a y value that makes the expression true. So if I were to draw that graph I would have this:  So lets just say I pick x = 0, y = -2. If I plug those into the equation I have 0 - 2 >= 2 Well, -2 is not >= 2, so I cant shade any area that contains the point (0,-2). But if I pick x = 1 and y = 3. I can plug those values into the equation and get: 1 + 3 >= 2 4 >= 2 Well of course 4 is greater than 2. That means I can shade the whole region that has the point (1,3) Basically youre checking coordinate points and seeing if they make the equation true: